OpenStudy (anonymous):

2/5y+6=10 how do get my answer

OpenStudy (anonymous):

subtract 6 to both sides and then multiply by the recipricol (5/2) and you will get your answer for y

OpenStudy (anonymous):

2/5y=4 y=4(5/2) y=10

OpenStudy (anonymous):

thank you got it. I wasn't sure if I was suppose to add 2/5 or multiply by recipricol 5/2.